Lysozome is an enzyme that breaks bacterial cells. A solution containing 150 mg lysozome in 210 mL of solution has an osmotic pressure of 0.953 torr at 25.0°C. Determine the molar mass of lysozome.


1
Expert's answer
2021-04-02T05:17:15-0400

Solution:

Osmotic pressure (Π) obeys a law:

Π = M × R × T

where M is the molarity of the solution, R is the ideal gas constant, and T is the absolute temperature.


Lysozome is a nonelectrolyte, so we can solve equation above for the molarity corresponding to this osmotic pressure: M = Π / (R × T)


Π = 0.953 torr × (0.001316 atm / 1 torr) = 0.001254 atm

R = 0.0821 atm L mol-1 K-1

T = 25.0°C + 273 = 298 K

Hence,

Molarity of lysozome solution = (0.001254 atm) / (0.0821 atm L mol-1 K-1 × 298 K) = 0.00005126 mol/L

Molarity of lysozome solution = 0.00005 mol/L


(0.15 g lysozome) × (1 L soln/ 0.00005126 mol lysozome) × (1 / 0.21 L soln) = 13934.56 g/mol

The molar mass of lysozome = 13934.56 g/mol = 1.40×104 g/mol.


Answer: The molar mass of lysozome is 1.40×104 g/mol.